Circular No. 8303

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004ak, 2004al, AND 2004ap T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Cen-

ter for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2004al (cf. IAUC 8297), obtained by P. Berlind on Mar. 13.27 UT with the Mt. Hopkins 1.5-m telescope (+ FAST), shows it to be a type-II supernova. A narrow Hα emission line superposed on the spectrum of the supernova indicates a recession velocity of 4340 km/s for the host galaxy. The supernova expansion velocity, derived from the minimum of the Hβ line, is 6300 km/s. A spectrum of SN 2004ap (cf. IAUC 8300), obtained by Berlind on Mar. 13.25, shows it to be a type-Ia supernova about a week past maximum light. A spectrum of SN 2004ak (cf. IAUC 8297), obtained on Mar. 14.31, shows it to be a type-II supernova; the spectrum consists of a very blue continuum with P-Cyg lines of hydrogen and helium. Adopting the NED recession velocity of 7214 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 10700 km/s.

IRAS 05436–0007 T. Simon, University of Hawaii; and S. D. Brittain, E. L. Gibb, and

T. W. Rettig, University of Notre Dame, report on a high-resolution Keck NIRSPEC spectrum of the near-infrared counterpart to IRAS 05436–0007 (cf. IAUC 8284). The 2–5-μm data (Feb. 27.3 UT) show a substantial amount of cold polar and apolar (cf. Ehrenfreund et al. 1997, A.Ap. 328,

649) CO ice. Initial model fits to the ice profiles are consistent with predominantly cold (< 30 K) amorphous water ice and cold, predominantly apolar CO ice (both have optical depth τ ∼ 0.6). Also present are broadened emission lines of ^{12}CO (1-0), (2-1), and (3-2), as well as ^{13}CO, which likely originate from warm gas (a few thousand degrees K) in an inner accretion disk region (< 1 AU). Narrow ^{12}CO absorption components are superposed on the underlying emission features. The narrow absorption components, with the same radial velocity as the emission lines, suggest substantial cold circumstellar material from a flared disk (cf. Brittain et al. 2004, Ap.J., submitted). In contrast to the findings published on IAUC 8301, the hydrogen lines are not strongly detected in the NIRSPEC data. The hot and cold gas surrounding the central star are typical of a young T-Tau star, but the relatively deep water and CO ice features together would be unique for the class. Both absorptions indicate that the underlying star is embedded deeply within the Orion L1630 cloud and thus at an earlier stage of evolution, very possibly a class-I object in transition.

2004 March 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green